OpenStudy (anonymous):

the perimeter of a rectangle with length, l and width, w is given by P=2w+2l. If the perimeter of a certain rectangle is 105 feet and the length is 35 feet, what is the width? Please show work.... thanks

OpenStudy (anonymous):

P=2w+2l P=105 l=35 105=2w+2(35) 105=2w+70 105-70=2w+70-70 35=2w 35/2=w